# Review: Quantum Metrology and Sensing with Many-Body Systems

### v2

@@ VI Equilibrium quantum sensing: Topological Phase Transitions
Second order quantum phase transitions are known for various features including spontaneous breaking of a continuous symmetry, local order parameters, gap closing and scale invariance behavior. While such critical phenomena is usually good for quantum-enhanced sensitivity, it is not clear which of these features are responsible for achieving quantum enhanced sensitivity. Moreover, although these models are not as noise-sensitive as interferometric schemes, detailed analysis shows reduction in the QFI scaling with increasing noise Chen et al. (2021). Thus, building quantum many-body sensors based on beyond second-order phase transitions is crucial. There also exists phase transitions where the corresponding order parameter is global. These are called topological phase transitions and are characterized by discrete jump in topological indices of the ground state of the physical system such as the Chern number (Zhang, 2008; Stanescu, 2016; Goldman et al., 2016). In addition, topological systems are expected to be naturally more robust against local noise and thus their performance as a sensor be less noise-affected. In this section, we consider two different topological sensors, namely Symmetry Protected Topological (SPT) systems and non-Hermitian topological probes.
@@ VI.1 Symmetry protected topological sensors
Topological phase transitions, exemplified through SPT transitions, are fundamentally different from conventional second-order quantum criticality. In fact, SPT transitions are characterized by: (i) the emergence of robust edge/surface states which are protected against symmetry-preserving local perturbations Alldridge et al. (2020); Alase (2019); (ii) an integer-valued nonlocal quantity called a topological invariant Bernevig (2013); and (iii) the absence of long-range correlations and entanglement Chen et al. (2013).
One of the simplest and earliest examples of many-body systems with SPT transition is the Su-Schrieffer-Heeger (SSH) model (Su et al., 1979) which consists of one excitation in a 1D tight-binding lattice consisting of two different sublattices characterized by bonds of alternating strengths. The Hamiltonian is given by
$H_{\rm SSH}=-\sum_{j\in L}\left(t_{1}b_{j}^{\dagger}a_{j}+t_{2}a_{j+1}^{\dagger}b_{j}\right).$ (58)
where $t_{1}$ and $t_{2}$ are intra- and inter-site hopping rates. The parameter of interest is $\lambda{=}t_{1}/t_{2}$. For 2D and higher dimensional systems, the quintessential examples of topological systems are the so called Chern insulators and generalizations thereof (see Ref. (Qi and Zhang, 2011) for a detailed review). For a simple system that has been realized in optical lattices with cold atoms (Wu et al., 2016b), the spin-orbit coupled Hamiltonian is given by
$\hat{H}^{\rm Ch}=\displaystyle\sum_{\bm{k}}\begin{bmatrix}\hat{c}^{{\dagger}}_{\bm{k},\uparrow}&\hat{c}^{{\dagger}}_{\bm{k},\downarrow}\end{bmatrix}\enspace H_{\bm{k}}^{\rm Ch}\enspace\begin{bmatrix}\hat{c}_{\bm{k},\uparrow}&\hat{c}_{\bm{k},\downarrow}\end{bmatrix}^{T}\,,$ (59)
@@
For topological systems, Ginzburg-Landau theory fails since local order parameters do not exist and no continuous symmetry is broken. Hence the fidelity susceptibility or QFI based approach to detect topological phase transitions holds special relevance for such models, in addition to the practical task of building many-body quantum sensors. Extensive studies in this regard has been made on lower dimensional systems, including 1D Kitaev chain (Mondal et al., 2024), and Kitaev’s deformed toric code models (Abasto et al., 2008) and honeycomb models (Yang et al., 2008; Zhao and Zhou, 2009) in 2D. Gu and Lin (Gu and Lin, 2009) also pointed out that the critical exponent of fidelity for the honeycomb model is super-extensive and different for two sides of the critical point. Thus, fidelity susceptibility is indeed an useful tool for detecting topological phase transitions. In (Sarkar et al., 2022), the possibility of quantum enhancement of sensing in edge and ground states of topologically non-trivial single-excitation subspaces of generic free-Fermion Hamiltonians was investigated. For edge states, $N^{\beta}$ scaling, with $\beta{=}2$ of QFI was obtained at the transition point and constant scaling away from criticality, as shown in the top panel of Fig. 7. Significantly, the optimal measurement for saturating the Cramér-Rao bound was revealed as a constant site-location measurement on the lattice for every value of the parameter $\theta$. For many-body ground states, quadratic scaling of QFI was again obtained around the topological phase transition point, with classical linear scaling away from criticality, as shown in the bottom panel of Fig. 7.
@@
Figure 7: QFI scaling near transition. Top panels : Scaling exponent of QFI of edge state (for $\lambda<\lambda_{\rm c}$) and corresponding bulk state (for $\lambda>\lambda_{\rm c}$) as a function of $\lambda$ for (a) SSH model, and (b) Chern insulator model. Bottom panels : Scaling exponent of QFI of many-body ground state as a function of $\lambda$ subject to PBC (blue) and OBC (red) for (a) SSH model, and (b) Chern insulator model. Figures taken from Ref. (Sarkar et al., 2022).
@@ VI.2 Non-Hermitian Quantum Sensors
Non-Hermitian Hamiltonians provide an effective description of open system dynamics Ashida et al. (2020). The topological systems bear many surprising features due to the unique properties of the eigenstates and complex eigenvalues Bergholtz et al. (2021). In the absence of degeneracy, the eigenstates are linearly independent and form a complete basis, but they are not orthogonal. This stems from the fact that the right- and left-eigenstates are not the same. For a non-Hermitian Hamiltonian $H_{\rm NH}$, they are defined as
$H_{\rm NH}\ket{\psi^{\rm R}_{n}}$ $=E_{n}\ket{\psi^{\rm R}_{n}},$
$\bra{\psi^{\rm L}_{n}}H_{\rm NH}$ $=E_{n}\bra{\psi^{\rm L}_{n}}\implies H_{\rm NH}^{{\dagger}}\ket{\psi^{\rm L}_{n}}=E^{*}_{n}\ket{\psi^{\rm L}_{n}},$ (60)
@@
where $E_{n}$ is the corresponding eigenvalue. However, these states are biorthogonal Brody (2013), and upon normalization gives, $\braket{\tilde{\psi}^{\rm L}_{m}}{\tilde{\psi}^{\rm R}_{n}}{=}\delta_{mn}$. Here, |$\ket{\tilde{\psi}^{\rm R}_{n}}{=}\ket{\psi^{\rm R}_{n}}/\sqrt{\braket{\psi^{\rm L}_{n}}{\psi^{\rm R}_{n}}}$, and |$\ket{\tilde{\psi}^{\rm L}_{n}}{=}\ket{\psi^{\rm L}_{n}}/\sqrt{\braket{\psi^{\rm L}_{n}}{\psi^{\rm R}_{n}}}$. This helps in writing down the completeness relation $\sum_{n}\ket{\tilde{\psi}^{\rm R}_{n}}\bra{\tilde{\psi}^{\rm L}_{n}}{=}\mathds{1}$ and the spectral decomposition $H_{\rm NH}{=}\sum_{n}E_{n}\ket{\tilde{\psi}^{\rm R}_{n}}\bra{\tilde{\psi}^{\rm L}_{n}}$ |. Non-Hermitian Hamiltonians can have degeneracies analogous to Hermitian systems, i.e. identical eigenvalues with distinct eigenstates. The points in the Hamiltonian’s parameter space where this happens are known as diabolic points. There are also parameter values where more than one eigenvalues and the corresponding eigenstates can become identical. These are known as Exceptional Points (EP) and have no Hermitian counterpart. The right- and left-eigenstates corresponding to the EP are orthogonal to each other and the biorthogonal framework breaks down as the Hamiltonian is not digonalizable anymore. If the Hamiltonian has parity-time (PT) symmetry, then it commutes with the antilinear and antiunitary $\mathcal{PT}$ operator. For example, in spin-$1{/}2$ systems on a lattice, the parity operator is $\mathcal{P}{=}\sum_{j}\ket{-j}\bra{j}\otimes\sigma^{x}_{j}$, and the time-reversal operator is $\mathcal{T}{=}\sum_{j}\ket{j}\bra{j}\otimes\sigma^{y}_{j}K_{j}$. Here the term $|{-j}\rangle\langle j|$ swaps the position of spins at sites $j$ and $-j$. The action of $K_{j}$ is complex conjugation, and $\sigma^{(x,y,z)}_{j}$ are the Pauli spin operators on $j$-th site that affect the spin degree of freedom Krasnok et al. (2021); Mochizuki et al. (2016). In an appropriate parameter regime, a PT-symmetric Hamiltonian supports a real energy spectrum, and it is known as the PT-unbroken phase. In this case, the Hamiltonian and the PT operator share the same eigenstates, known as PT-symmetric eigenstates. Upon variation of the Hamiltonian parameters, a PT-broken phase can be entered where some or all eigenvalues become complex that appear in complex conjugate pairs. This happens through a phase transition associated with spontaneous PT-symmetry breaking. The PT-symmetry breaking point is an EP of the system.
Figure 8: EP based and non-reciprocity based sensors. (a) Energy slitting versus perturbation strength $\epsilon$ near EP shows enhancement over standard splitting near a diabolic point (DP). For EP of order 2 in Ref. Chen et al. (2017), the black line shows the expected $\epsilon^{1/2}$ behavior. The inset shows scaling exponent of EP based sensor (blue squares) increasing form $1/2$ (black line) as $\epsilon$ increases. The exponent for DP based sensor is $1$ (red circles). Figure adapted from Ref. Chen et al. (2017). (b) For a PT-symmetric system in Ref. Yu et al. (2020), EP of order 2 appears at the symmetry breaking point. The black circles are the experimental data for the energy splitting with perturbation $\epsilon$. The red line shows the expected $\sqrt{|\epsilon|}$ behavior. In the PT-unbroken regime, the energies are real and the upper panel shows the splitting. In the PT-broken regime of the system considered , the real part of energy stays constant while the splitting is observed in the imaginary part (lower panel). As the EP is approached, deviation from the expected value occurs due to excessive noise. Figure adapted from Ref. Yu et al. (2020). (c) The signal-to-noise ratio (SNR) shows the exponentially enhanced sensitivity with system size $N$ in the non-reciprocity based sensing protocol in Ref. McDonald and Clerk (2020). It is based on the Hatano-Nelson model in Eq. (61) and $A=\log{(J_{R}/J_{L})}/2$. The analysis is based on linear response theory which breaks down for very large $N$, due to amplification of noise. The expected value of $N$ for this to happen with fixed values of other parameters is shown with dashed vertical line. Figure adapted from Ref. McDonald and Clerk (2020). (d) QFI ($\mathcal{I}$) with respect to perturbation coupling the boundaries ($\Gamma$) as a function of system size $N$ for the non-Hermitian SSH model based proposal in Ref. Koch and Budich (2022). Here $t_{1}$ and $t_{2}$ are the intra- and inter-site hopping terms. The QFI increases exponentially with coefficient $\alpha$ (solid line) until it saturates to a value that increases with decreasing $\Gamma$. Figure adapted from Ref. Koch and Budich (2022).
@@
EPs are branch point singularities in the complex energy manifold and generate non-trivial topology, quantified by winding of eigenvalues and eigenvectors Leykam et al. (2017). Sensing based on EP was proposed first in Ref. Wiersig (2014) for coupled cavity systems – with photon loss and gain terms – that have resonant modes at EP. A non-Hermitian perturbation with strength $\epsilon$ would move the system away from the EP. For an EP of degree $n$ (i.e. $n$ eigenstates coalescing), the resulting energy splitting ${\sim}\epsilon^{1/n}$. In contrast, for Hermitian degeneracies or diabolic points, the leading order splitting is ${\sim}\epsilon$. The advantage of the EP based sensors can be quantified by the susceptibility $d_{\epsilon}\epsilon^{1/n}\propto\epsilon^{-(1-1/n)}$, which diverges as $\epsilon\to 0$. This was experimentally demonstrated with optical microring resonators Chen et al. (2017); Hodaei et al. (2017) and also with further imposition of PT symmetry Liu et al. (2016); Yu et al. (2020). However, these schemes do not reveal the precision of the sensor which is the smallest change in $\epsilon$ that can be measured. The standard measurement is done by first implementing an excitation, followed by analyzing the resonant scattering output. The signal-to-noise ratio of the output is a measure of QFI, which sets the precision. As the eigenstates participating in the excitation-deexcitation process become more and more non-orthogonal while approaching EP, it was debated that the resulting quantum noise would overwhelm the advantage of EP based sensors Langbein (2018); Chen et al. (2019). Experimental observation also confirmed this Wang et al. (2020). Signal enhancement was nevertheless predicted at specific frequencies Zhang et al. (2019) or with additional gain terms Lau and Clerk (2018) and was also observed experimentally for systems slightly detuned from EP Kononchuk et al. (2022). The effectiveness of EP based sensors is still a topic of active research Wiersig (2020).
@@
$H_{\rm HN}=\sum_{j}\left(J_{L}\ket{j}\bra{j{+}1}+J_{R}\ket{j{+}1}\bra{j}\right).$ (61)
The imbalanced hopping terms generate an unique non-Hermitian topology even in the single-band structure, unlike Hermitian cases. Considering a $N$-site chain of coupled bosonic modes with a classical resonant drive at site 1, it was shown that the dynamics of the two quadrature modes are governed by a ‘doubled’ Hatano-Nelson Hamiltonian, with opposite reciprocities for the two modes. This imposes a $\mathbb{Z}_{2}$ symmetry which is broken by a perturbation on site $N$ that mixes the two modes with strength $\epsilon$. An input drive in terms of one mode would generate a large signal as the output in the other mode in the steady state. With judicious choice of the optimal homodyne measurement Banchi et al. (2015), the signal-to-noise ratio gives the QFI with respect to $\epsilon$ which scales exponentially with system size, namely QFI $\sim|\epsilon|e^{(N-1)\log{(J_{R}/J_{L})}/2}$. This kind of exponential sensitivity was also reported in Ref. Koch and Budich (2022) with a similar coupled cavity system with input drive but here the perturbation term – with strength $\epsilon$ – connects the boundary modes (sites 1 and $N$). Here, the dynamics is governed by a non-Hermitian extension of the SSH model with balanced loss and gain terms on alternating sites. This breaks the sublattice symmetry, but makes the systems PT-symmetric. In the topologically non-trivial regime, the signal-to-noise ratio of the steady state output field scales exponentially with system size, i.e. QFI $\sim e^{\alpha N}$, where $\alpha$ is a positive coefficient. This behavior is closely connected to the acute spectral sensitivity of non-Hermitian topological systems Koch and Budich (2020); Kunst et al. (2018); Edvardsson and Ardonne (2022). In particular, exponential sensitiveness of the edge state energy $E_{0}$ to the coupling at the boundary $\epsilon$ (i.e. $\partial_{\epsilon}E_{0}\sim e^{\alpha N}$), has also led to sensing proposal in Ref. Budich and Bergholtz (2020). These proposals function anywhere in the topologically non-trivial regime and does not require fine-tuning to an EP. An experimental confirmation of edge state sensitivity at the EP was also performed recently with an optical resonator setup Guo et al. (2021b). Even in the absence of EP, superiority of sensing has been shown experimentally Xiao et al. (2024) and Heisenberg scaling in time has also been demonstrated with non-Unitary dynamics Yu et al. (2024b). Such scaling behaviors are quite promising regarding sensing in open systems, although recent theoretical arguments cast doubts about actual advantages of non-Hermitian sensors over their Hermitian counterparts Ding et al. (2023).
Figure 10: Bulk Hamiltonian parameter estimation with Hatano-Nelson model with $\lambda=J_{R}/J_{L}$ in Eq. (61). (a) Bulk spectrum with periodic boundary conditions (PBC). (b) Cumulative site-population for open boundary conditions (OBC) eigenstates shows the non-Hermitian skin effect and its absence (inset). (c) Quadratic scaling of QFI with system size $L$ at the transition $\lambda=\lambda_{c}=\pm 1$. (d) QFI Scaling exponent near transition. Figures taken from Ref. Sarkar et al. (2024).
@@
Beyond sensing boundary perturbations, non-Hermitian topology also can be used for estimating bulk Hamiltonian parameters. The notion of band gaps are extended in the non-Hermitian regime to two different gap structures: line gap and point gap (see Fig. 9Kawabata et al. (2019). A spectral loop (for 1D systems) or finite spectral area (in higher dimensions) correspond to the presence of non-Hermitian skin effect in finite systems Martinez Alvarez et al. (2018); Yao and Wang (2018), where most of the eigenstates are localized at a boundary Borgnia et al. (2020); Zhang et al. (2020, 2022); Li et al. (2020). Changes in the bulk Hamiltonian parameters can make the spectral structure contract to a curve with zero spectral area. This signals a closure of the point gap, which makes the skin effect vanish. The corresponding localization transition in the eigenstates can be used for sensing, as was shown in Ref. Sarkar et al. (2024), to achieve precision with Heisenberg scaling. This extends the role of gap closing for quantum-enhanced sensitivity to a truly non-Hermitian domain. For the aforementioned Hatano-Nelson model, this scaling can be obtained analytically (see Fig. 10). The robustness against local disorders is also shown by mapping it to a non-Hermitian Aubrey-André-Harper model Longhi (2019). Apart form extension to prototypical multi-band models Yao and Wang (2018) and 2D systems Yao et al. (2018), experimental implementation is also possible in discrete-time non-Unitary quantum walk setups Xiao et al. (2020, 2021).

### v3

@@ VI Equilibrium quantum sensing: Topological Phase Transitions
Second order quantum phase transitions are known for various features including spontaneous breaking of a continuous symmetry, local order parameters, gap closing and scale invariance behavior. While such critical phenomena is usually good for quantum-enhanced sensitivity, it is not clear which of these features are responsible for achieving quantum-enhanced sensitivity. Moreover, although these models are not as noise-sensitive as interferometric schemes, detailed analysis shows reduction in the QFI scaling with increasing noise Chen et al. (2021). Thus, building quantum many-body sensors based on beyond second-order phase transitions is crucial. There also exists phase transitions where the corresponding order parameter is global. These are called topological phase transitions and are characterized by discrete jump in topological indices of the ground state of the physical system such as the Chern number (Zhang, 2008; Stanescu, 2016; Goldman et al., 2016). In addition, topological systems are expected to be naturally more robust against local noise and thus their performance as a sensor be less noise-affected. In this section, we consider two different topological sensors, namely Symmetry Protected Topological (SPT) systems and non-Hermitian topological probes.
@@ VI.1 Symmetry protected topological sensors
Topological phase transitions, exemplified through SPT transitions in free-fermionic systems, are fundamentally different from conventional second-order quantum criticality. In fact, SPT transitions are characterized by: (i) the emergence of robust edge/surface states which are protected against symmetry-preserving local perturbations Alldridge et al. (2020); Alase (2019); (ii) an integer-valued nonlocal quantity called a topological invariant Bernevig (2013); and (iii) the absence of long-range correlations and entanglement Chen et al. (2013).
One of the simplest and earliest examples of free-fermionic systems with SPT transition is the Su-Schrieffer-Heeger (SSH) model (Su et al., 1979) which consists of one excitation in a 1D tight-binding lattice with $N$ sites consisting of two different sublattices characterized by bonds of alternating strengths. The Hamiltonian is given by
$\hat{H}_{\rm SSH}=-\sum_{j}\left(t_{1}\hat{b}_{j}^{\dagger}\hat{a}_{j}+t_{2}\hat{a}_{j+1}^{\dagger}\hat{b}_{j}\right).$ (58)
where $t_{1}$ and $t_{2}$ are intra- and inter-site hopping rates and $\hat{a}_{j}$, $\hat{b}_{j}$ are the annihilation operators for the two sublattices. The parameter of interest is $\theta{=}t_{1}/t_{2}$. For 2D and higher dimensional systems, the quintessential examples of topological systems are the so called Chern insulators and generalizations thereof (see Ref. (Qi and Zhang, 2011) for a detailed review). For a simple system that has been realized in optical lattices with cold atoms (Wu et al., 2016b), the spin-orbit coupled Hamiltonian is given by
$\hat{H}^{\rm Ch}=\displaystyle\sum_{\bm{k}}\begin{bmatrix}\hat{c}^{{\dagger}}_{\bm{k},\uparrow}&\hat{c}^{{\dagger}}_{\bm{k},\downarrow}\end{bmatrix}\enspace\hat{H}_{\bm{k}}^{\rm Ch}\enspace\begin{bmatrix}\hat{c}_{\bm{k},\uparrow}&\hat{c}_{\bm{k},\downarrow}\end{bmatrix}^{T}\,,$ (59)
@@
For topological systems, Ginzburg-Landau theory fails since local order parameters do not exist and no continuous symmetry is broken. Hence the fidelity susceptibility or QFI based approach to detect topological phase transitions holds special relevance for such models, in addition to the practical task of building many-body quantum sensors. Extensive studies in this regard has been made on lower dimensional systems, including 1D Kitaev chain (Mondal et al., 2024), and Kitaev’s deformed toric code models (Abasto et al., 2008) and honeycomb models (Yang et al., 2008; Zhao and Zhou, 2009) in 2D. Gu and Lin (Gu and Lin, 2009) also pointed out that the critical exponent of fidelity for the honeycomb model is super-extensive and different for two sides of the critical point. Thus, fidelity susceptibility is indeed an useful tool for detecting topological phase transitions. In (Sarkar et al., 2022), the possibility of quantum enhancement of sensing in edge and ground states of topologically non-trivial single-excitation subspaces of generic free-Fermion Hamiltonians was investigated. For edge states, $N^{\beta}$ scaling, with $\beta{=}2$ of QFI was obtained at the transition point and constant scaling away from criticality, as shown in Figs. 7(a) and (b). Significantly, the optimal measurement for saturating the Cramér-Rao bound was revealed as a constant site-location measurement on the lattice for every value of the parameter $\theta$. For many-body ground states, quadratic scaling of QFI was again obtained around the topological phase transition point, with classical linear scaling away from criticality, as shown in Figs. 7(c) and (d).
@@
Figure 7: QFI scaling near transition. Top panels : Scaling exponent of QFI of edge state (for $\theta<\theta_{\rm c}$) and corresponding bulk state (for $\theta>\theta_{\rm c}$) as a function of $\theta$ for (a) SSH model, and (b) Chern insulator model. Bottom panels : Scaling exponent of QFI of many-body ground state as a function of $\theta$ subject to PBC (blue) and OBC (red) for (c) SSH model, and (d) Chern insulator model. Figures adapted from Ref. (Sarkar et al., 2022).
@@ VI.2 Non-Hermitian quantum sensors
Non-Hermitian Hamiltonians can generate non-Unitary dynamics which is closely connected with open system dynamics Ashida et al. (2020). For example, in the Lindbladian description, the evolution of the density operator can be expressed in terms of the properties of the non-Hermitian Liouvillian super-operator. Also, for the dynamics during which no Lindbladian jump operator acts (the probability of which decreases exponentially in time), an effective non-Hermitian Hamiltonian can describe a pure state evolution. With recent experimental progress, it is possible to design specific non-Hermitian Hamiltonians that determine the evolution of the system. In particular, the topological systems bear many surprising features due to the unique properties of the eigenstates and complex eigenvalues Bergholtz et al. (2021). In the absence of degeneracy, the eigenstates are linearly independent and form a complete basis, but they are not orthogonal. This stems from the fact that the right- and left-eigenstates are not the same. For a non-Hermitian Hamiltonian $\hat{H}_{\rm NH}$, they are defined as
$\hat{H}_{\rm NH}\ket{\psi^{\rm R}_{n}}$ $=E_{n}\ket{\psi^{\rm R}_{n}},$
$\bra{\psi^{\rm L}_{n}}\hat{H}_{\rm NH}$ $=E_{n}\bra{\psi^{\rm L}_{n}}\implies\hat{H}_{\rm NH}^{{\dagger}}\ket{\psi^{\rm L}_{n}}=E^{*}_{n}\ket{\psi^{\rm L}_{n}},$ (60)
[…]
@@
Non-Hermitian Hamiltonians can have degeneracies analogous to Hermitian systems, i.e. identical eigenvalues with distinct eigenstates. The points in the Hamiltonian’s parameter space where this happens are known as diabolic points. There are also parameter values where more than one eigenvalues and the corresponding eigenstates can become identical. These are known as Exceptional Points (EP) and have no Hermitian counterpart. The right- and left-eigenstates corresponding to the EP are orthogonal to each other and the biorthogonal framework breaks down as the Hamiltonian is not digonalizable anymore. If the Hamiltonian has parity-time (PT) symmetry, then it commutes with the antilinear and antiunitary $\mathcal{PT}$ operator. For example, in spin-$1{/}2$ systems on a lattice, the parity operator is $\hat{\mathcal{P}}{=}\sum_{j}\ket{-j}\bra{j}\otimes\hat{\sigma}^{x}_{j}$, and the time-reversal operator is $\hat{\mathcal{T}}{=}\sum_{j}\ket{j}\bra{j}\otimes\hat{\sigma}^{y}_{j}\hat{K}_{j}$. Here the term $|{-j}\rangle\langle j|$ swaps the position of spins at sites $j$ and $-j$. The action of $\hat{K}_{j}$ is complex conjugation, and $\hat{\sigma}^{(x,y,z)}_{j}$ are the Pauli spin operators on $j$-th site that affect the spin degree of freedom Krasnok et al. (2021); Mochizuki et al. (2016). In an appropriate parameter regime, a PT-symmetric Hamiltonian supports a real energy spectrum, and it is known as the PT-unbroken phase. In this case, the Hamiltonian and the PT operator share the same eigenstates, known as PT-symmetric eigenstates. Upon variation of the Hamiltonian parameters, a PT-broken phase can be entered where some or all eigenvalues become complex that appear in complex conjugate pairs. This happens through a phase transition associated with spontaneous PT-symmetry breaking. The PT-symmetry breaking point is an EP of the system.
@@
EPs are branch point singularities in the complex energy manifold and generate non-trivial topology, quantified by winding of eigenvalues and eigenvectors Leykam et al. (2017). Sensing based on EP was proposed first in Ref. Wiersig (2014) for coupled cavity systems – with photon loss and gain terms – that have resonant modes at EP. A non-Hermitian perturbation with strength $\epsilon$ would move the system away from the EP. For an EP of degree $n$ (i.e. $n$ eigenstates coalescing), the resulting energy splitting ${\sim}\epsilon^{1/n}$. In contrast, for Hermitian degeneracies or diabolic points, the leading order splitting is ${\sim}\epsilon$. The advantage of the EP based sensors can be quantified by the susceptibility $d_{\epsilon}\epsilon^{1/n}\propto\epsilon^{-(1-1/n)}$, which diverges as $\epsilon\to 0$. This was experimentally demonstrated with optical microring resonators Chen et al. (2017); Hodaei et al. (2017) and also with further imposition of PT symmetry Liu et al. (2016); Yu et al. (2020) (see Fig. 8). However, these schemes do not reveal the precision of the sensor which is the smallest change in $\epsilon$ that can be measured. The standard measurement is done by first implementing an excitation, followed by analyzing the resonant scattering output. The signal-to-noise ratio of the output is a measure of the precision, which corresponds to the QFI if the associated observable can be chosen to be optimal. As the eigenstates participating in the excitation-deexcitation process become more and more non-orthogonal while approaching EP, it was debated that the resulting quantum noise would overwhelm the advantage of EP based sensors Langbein (2018); Chen et al. (2019). Experimental observation also confirmed this Wang et al. (2020). Signal enhancement was nevertheless predicted at specific frequencies Zhang et al. (2019) or with additional gain terms Lau and Clerk (2018) and was also observed experimentally for systems slightly detuned from EP Kononchuk et al. (2022). The effectiveness of EP based sensors is still a topic of active research Wiersig (2020).
Figure 8: EP based and non-reciprocity based sensors. (a) Energy slitting versus perturbation strength $\epsilon$ near EP shows enhancement over standard splitting near a diabolic point (DP). For EP of order 2 in Ref. Chen et al. (2017), the black line shows the expected $\epsilon^{1/2}$ behavior. The inset shows scaling exponent of EP based sensor (blue squares) increasing form $1/2$ (black line) as $\epsilon$ increases. The exponent for DP based sensor is $1$ (red circles). Figure adapted from Ref. Chen et al. (2017). (b) For a PT-symmetric system in Ref. Yu et al. (2020), EP of order 2 appears at the symmetry breaking point. The black circles are the experimental data for the energy splitting with perturbation $\epsilon$. The red line shows the expected $\sqrt{|\epsilon|}$ behavior. In the PT-unbroken regime, the energies are real and the upper panel shows the splitting. In the PT-broken regime of the system considered , the real part of energy stays constant while the splitting is observed in the imaginary part (lower panel). As the EP is approached, deviation from the expected value occurs due to excessive noise. Figure adapted from Ref. Yu et al. (2020). (c) The signal-to-noise ratio (SNR) shows the exponentially enhanced sensitivity with system size $N$ in the non-reciprocity based sensing protocol in Ref. McDonald and Clerk (2020). It is based on the Hatano-Nelson model in Eq. (61) and $A=\log{(J_{R}/J_{L})}/2$. The analysis is based on linear response theory which breaks down for very large $N$, due to amplification of noise. The expected value of $N$ for this to happen with fixed values of other parameters is shown with dashed vertical line. Figure adapted from Ref. McDonald and Clerk (2020). (d) CFI with respect to perturbation coupling the boundaries ($\Gamma$) as a function of system size $N$ for the non-Hermitian SSH model based proposal in Ref. Koch and Budich (2022). Here $t_{1}$ and $t_{2}$ are the intra- and inter-site hopping terms. The CFI increases exponentially with coefficient $\alpha$ (solid line) until it saturates to a value that increases with decreasing $\Gamma$. Figure adapted from Ref. Koch and Budich (2022).
$\hat{H}_{\rm HN}=\sum_{j}\left(J_{L}\ket{j}\bra{j{+}1}+J_{R}\ket{j{+}1}\bra{j}\right).$ (61)
The imbalanced hopping terms generate an unique non-Hermitian topology even in the single-band structure, unlike Hermitian cases. Considering a $N$-site chain of coupled bosonic modes with a classical resonant drive at site 1, it was shown that the dynamics of the two quadrature modes are governed by a ‘doubled’ Hatano-Nelson Hamiltonian, with opposite reciprocities for the two modes. This imposes a $\mathbb{Z}_{2}$ symmetry which is broken by a perturbation on site $N$ that mixes the two modes with strength $\epsilon$. An input drive in terms of one mode would generate a large signal as the output in the other mode in the steady state. With judicious choice of the optimal homodyne measurement Banchi et al. (2015), the signal-to-noise ratio gives the QFI with respect to $\epsilon$ which scales exponentially with system size, namely QFI $\sim|\epsilon|e^{(N-1)\log{(J_{R}/J_{L})}/2}$. This kind of exponential sensitivity was also reported in Ref. Koch and Budich (2022) with a similar coupled cavity system with input drive but here the perturbation term – with strength $\epsilon$ – connects the boundary modes (sites 1 and $N$). Here, the dynamics is governed by a non-Hermitian extension of the SSH model with balanced loss and gain terms on alternating sites. This breaks the sublattice symmetry, but makes the systems PT-symmetric. In the topologically non-trivial regime, the sensitivity of the steady state output field scales exponentially with system size, i.e. ${\sim}e^{\alpha N}$, where $\alpha$ is a positive coefficient. This behavior is closely connected to the acute spectral sensitivity of non-Hermitian topological systems Koch and Budich (2020); Kunst et al. (2018); Edvardsson and Ardonne (2022). In particular, exponential sensitiveness of the edge state energy $E_{0}$ to the coupling at the boundary $\epsilon$ (i.e. $\partial_{\epsilon}E_{0}\sim e^{\alpha N}$), has also led to sensing proposal in Ref. Budich and Bergholtz (2020). These proposals function anywhere in the topologically non-trivial regime and does not require fine-tuning to an EP. An experimental confirmation of edge state sensitivity at the EP was also performed recently with an optical resonator setup Guo et al. (2021b). Even in the absence of EP, superiority of sensing has been shown experimentally Xiao et al. (2024) and Heisenberg scaling in time has also been demonstrated with non-Unitary dynamics Yu et al. (2024b). Such scaling behaviors are quite promising regarding sensing in open systems, although recent theoretical arguments cast doubts about actual advantages of non-Hermitian sensors over their Hermitian counterparts Ding et al. (2023).
@@
Beyond sensing boundary perturbations, non-Hermitian topology also can be used for estimating bulk Hamiltonian parameters. The notion of band gaps are extended in the non-Hermitian regime to two different gap structures: line gap and point gap (see Fig. 9Kawabata et al. (2019). A spectral loop (for 1D systems) or finite spectral area (in higher dimensions) correspond to the presence of non-Hermitian skin effect in finite systems Martinez Alvarez et al. (2018); Yao and Wang (2018), where most of the eigenstates are localized at a boundary Borgnia et al. (2020); Zhang et al. (2020, 2022c); Li et al. (2020). Changes in the bulk Hamiltonian parameters can make the spectral structure contract to a curve with zero spectral area. This signals a closure of the point gap, which makes the skin effect vanish. The corresponding localization transition in the eigenstates can be used for sensing. As was shown in Ref. Sarkar et al. (2024), precision with Heisenberg scaling can be achieved using such probes. This extends the role of gap closing for quantum-enhanced sensitivity to a truly non-Hermitian domain. For the aforementioned Hatano-Nelson model, this scaling can be obtained analytically (see Fig. 10). The robustness against local disorders is also shown by mapping it to a non-Hermitian Aubrey-André-Harper model Longhi (2019). Apart form extension to prototypical multi-band models Yao and Wang (2018) and 2D systems Yao et al. (2018), experimental implementation is also possible in discrete-time non-Unitary quantum walk setups Xiao et al. (2020, 2021).
Figure 10: Bulk Hamiltonian parameter estimation with Hatano-Nelson model with $\theta=J_{R}/J_{L}$ in Eq. (61). (a) Bulk spectrum with periodic boundary conditions (PBC). (b) Cumulative site-population for open boundary conditions (OBC) eigenstates shows the non-Hermitian skin effect and its absence (inset). (c) Quadratic scaling of QFI with system size $N$ at the transition $\theta=\theta_{c}=\pm 1$. (d) QFI Scaling exponent near transition. Figures adapted from Ref. Sarkar et al. (2024).